# Linear spectroscopy of collective modes and the gap structure in two-dimensional superconductors

### v1

@@ Abstract
To conclude, we have shown that linear spectroscopy can be an invaluable tool for studying the pairing mechanisms in 2D superconductors. Using a simple but powerful toy model, we showed how various exotic collective modes should be detectable in linear response, either at $\mathbf{q}=0$ (in the case of the Leggett mode), or at finite $\mathbf{q}$ (in the case of the clapping modes). By considering the example of superconducting rhombohedral trilayer graphene (RTG) under a background vertical displacement field, we further showed how the observation of an in-gap clapping mode could yield compelling evidence for unconventional, all-electronic pairing mechanisms. We also showed how linear spectroscopy above the quasiparticle excitation gap reveals information on the gap structure, in the form of sharp features at frequencies corresponding to gap extrema over the Fermi surface.
The physics we have described should be accessible in experiments. The superconducting phases in RTG have transition temperatures at the scale of between 50 and 100 mK [27], so one expects the gap and possible in-gap clapping modes to reside at the scale of a few GHz to 10s of GHz, easily within the range of microwave function generators. Further, we expect that similar phenomenology may occur in other quasi-two-dimensional superconductors which are sensitive to a vertical displacement field, such as Bernal bilayer [39, 40] and twisted trilayer [41, 42, 43, 44] graphene. Future work should study the collective mode spectroscopy of these systems.
Acknowledgements: B. A. L. was hosted at the Institute for Theoretical Physics at the University of Cologne during a considerable portion of the preparation of this manuscript, and gratefully acknowledges their generous support, as well as the financial support of the Zuckerman STEM Leadership Program. This work was supported by NSF-BSF award DMR-2310312, by the European Union’s Horizon 2020 research and innovation programme (grant agreements LEGOTOP No. 788715 and HQMAT No. 817799), and the DFG CRC SFB/TRR183. M. R. gratefully acknowledges support from the Brown Investigator Award, a program of the Brown Science Foundation, the University of Washington College of Arts and Sciences, and the Kenneth K. Young Memorial Professorship.

### v2

@@ Abstract
To conclude, we have shown that simple linear spectroscopy via gate electrodes can be an invaluable tool for studying the pairing mechanisms in low-dimensional superconductors. Using a simple but powerful toy model, we showed how various exotic collective modes should be detectable in AC capacitance measurements. By considering the example of superconducting rhombohedral trilayer graphene (RTG) under a perpendicular displacement field, we further showed how the observation of an in-gap clapping mode could yield compelling evidence for unconventional, all-electronic pairing mechanisms. We also showed how linear spectroscopy above the quasiparticle excitation gap reveals information on the gap structure, in the form of sharp features at frequencies corresponding to gap extrema over the Fermi surface.
The physics we have described should be accessible in experiments. The superconducting phases of RTG [50] require dual gating to be realized, so RTG is an ideal candidate material for the gate-based detection scheme we have proposed. These phases have transition temperatures between $\sim 40$ and $\sim 100$ mK, so one expects the gap and possible in-gap clapping modes to reside at the scale of a few GHz to 10s of GHz, easily within the range of microwave function generators. Further, we expect that similar phenomenology may occur in other quasi-two-dimensional superconductors which are sensitive to a displacement field, such as Bernal bilayer [63, 64] and twisted trilayer [65, 66, 67, 68] graphene, as well as twisted bilayer WSe2 [69, 70]. Future work should study the collective mode spectroscopy of these systems.
Acknowledgements.
We thank Peter Hirschfeld, Saurabh Maiti, and Surajit Sarkar for their helpful comments on this manuscript. B. A. L. was hosted at the Institute for Theoretical Physics at the University of Cologne during a considerable portion of the preparation of this manuscript, and gratefully acknowledges their generous support, as well as the financial support of the Zuckerman STEM Leadership Program. This work was supported by NSF-BSF award DMR-2310312, by the European Union’s Horizon 2020 research and innovation programme (grant agreements LEGOTOP No. 788715 and HQMAT No. 817799), and the DFG CRC SFB/TRR183. M. R. gratefully acknowledges support from the Brown Investigator Award, a program of the Brown Science Foundation, the University of Washington College of Arts and Sciences, and the Kenneth K. Young Memorial Professorship.